Session: Oct 2008 (session 3)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Bo Xia at Tue Nov 4 19:33:21 2008) Success :

Medicina (orfei at Tue Nov 4 20:01:51 2008) Success :

Onsala (Michael Lindvist at Tue Nov 4 21:28:06 2008) Success :
No known problems.

Yebes (JF Desmurs at Wed Nov 5 16:06:09 2008) Minor failures : RECO SCHED
At beginning of the session, we get an access error message to the disk, we switched to the second one but lost the first scan (later P. de Vicente found and fixed the problem). Data lost between (UT) 11h53-12h14 (restart at line 914): Because of the problem detected during N08K4, we stopped observing the schedule to checked the IF setting by observing a strong H2O maser, every thing was OK. Few slewing warning messages but in most cases, the antenna was on source when starting writing the data (or delayed just by a couple of seconds).

Effelsberg (Uwe Bach at Thu Nov 6 11:00:41 2008) Success :

Noto (P. Cassaro at Tue Nov 11 22:13:14 2008) Success :